# SAC: Neural Speech Codec with Semantic-Acoustic Dual-Stream Quantization

### v1

@@ Abstract
Speech codecs that convert continuous speech signals into discrete tokens have become essential for speech language models (SLMs). However, existing codecs struggle to balance high-quality reconstruction with semantically rich representations, limiting their effectiveness in both generative and understanding tasks. In this work, we propose SAC, a neural speech codec with semantic–acoustic dual-stream quantization. By disentangling semantic and acoustic modeling into two dedicated streams, SAC enables each to be optimized for its respective role. Comprehensive evaluations show that SAC achieves strong reconstruction performance across diverse bitrates under both clean and noisy conditions, with particularly high scores on UTMOS and WER, demonstrating superior perceptual quality and intelligibility. Moreover, SAC substantially outperforms state-of-the-art codecs in semantic representation, achieving a level comparable to that of self-supervised learning (SSL) continuous embeddings. Finally, our analysis of speech disentanglement highlights the effectiveness of the dual-stream design, offering new potential for controllable speech applications. The code and pre-trained models are available at https://github.com/Soul-AILab/SAC.  [^1] Demo at https://sac-codec.github.io.
@@ 1 Introduction
With the rapid advancement of large language models (LLMs), speech language models (SLMs) have emerged by extending text-based LLMs with speech modalities (Peng et al., 2024; Ji et al., 2024a). Central to these models is the speech tokenizer, which discretizes continuous speech waveforms into token sequences, thereby enabling seamless integration with token-based language models (Guo et al., 2025). Leveraging such speech tokens, SLMs have driven progress in a wide range of downstream applications, including text-to-speech (TTS) (Xie et al., 2025; Wang et al., 2025b), speech understanding (Wang et al., 2024), and spoken dialogue systems (Xu et al., 2025; Zeng et al., 2024a; Chen et al., 2024b; Défossez et al., 2024).
Figure 1: Comparison of codecs on speech reconstruction. The x-axis shows WER, reflecting speech intelligibility, while the y-axis presents UTMOS, reflecting perceptual acoustic quality. Circle size indicates bitrate.
Semantic tokens are among the most widely used tokens in speech processing. These tokens are typically derived from either self-supervised models (Yang et al., 2024) or supervised models (Du et al., 2024a, b), making them effective in capturing semantic meaning. However, the absence of essential acoustic information significantly limits the applicability of semantic tokens. In contrast, acoustic tokens are usually generated by neural audio codecs trained with reconstruction objectives (Défossez et al., 2022; Zeghidour et al., 2021). While this approach preserves fine-grained acoustic details, the lack of semantic supervision results in weaker alignment with semantic content and reduced compatibility with text-based LMs (Wang et al., 2025a; Deng et al., 2025).
To enhance semantic representation, recent advancements in speech codecs have explored incorporating semantic supervision during training. For example, SpeechTokenizer (Zhang et al., 2023) employs semantic distillation, where representations from an SSL model guide the output of the first residual vector quantization (RVQ) layer. The X-Codec series (Ye et al., 2025a, b) and XY-Tokenizer (Gong et al., 2025) adopt an “X-shaped” paradigm, explicitly injecting pre-trained semantic features and fusing them with acoustic embeddings before quantization. However, while these approaches improve semantic alignment compared to methods without semantic constraints, they still fall short of pure semantic tokens in terms of semantic relevance. This raises a central question: Can semantic and acoustic tokens be disentangled at the token level, allowing each to specialize in its respective role? In this work, we explore such a dual-stream design that decouples semantic and acoustic modeling into independent streams, and show that it leads to improved semantic representation and reconstruction performance.
@@
Acoustic Dual-Stream Neural Speech Codec. Unlike prior approaches that inject semantic supervision into codecs, SAC complements the semantic tokens by introducing a separate acoustic token stream, which provides the essential acoustic information missing from the semantic tokens, all while ensuring the integrity of the semantic representations. Specifically, in the semantic stream, we adopt a pre-trained speech tokenizer (Zeng et al., 2024b) to extract semantic tokens aligned with linguistic content, keeping it frozen during training to ensure faithful retention of semantic information. In the acoustic stream, we follow the design of neural audio codecs (Kumar et al., 2023), employing temporally distributed acoustic tokens to capture the essential acoustic information, e.g., timbre and emotional attributes, that is missing from the semantic tokens. This dual-stream design unifies the complementary strengths of both within a single framework: semantic tokenizers excel in speech understanding and dialogue tasks (Zeng et al., 2024a; Ding et al., 2025), while acoustic tokenizers are particularly effective in generative modeling (Wang et al., 2023; Chen et al., 2024a). At the decoding stage, a ConvNeXt-based (Liu et al., 2022) prenet is employed to fuse the two streams of embeddings, followed by a codec decoder that reconstructs the waveform. Experiments show that SAC demonstrates solid performance in both reconstruction and semantic relevance. Our main contributions can be summarized as follows:
  • We propose SAC, a semantic–acoustic dual-stream neural speech codec. By explicitly disentangling speech encoding into parallel semantic and acoustic streams, SAC allows each stream to specialize in modeling linguistic content and acoustic detail, respectively.
  • SAC outperforms existing codecs with strong speech reconstruction and semantic representation across different bitrates. It attains superior scores in UTMOS and WER, reflecting excellent perceptual quality and intelligibility, while its semantic representation capability is even comparable to that of SSL models.
@@ 2 Related Work
Semantic tokens are derived from quantized representations of self-supervised learning (SSL) models or supervised models. They are strongly correlated with textual semantics but are not primarily designed for waveform reconstruction. Following the introduction of the $\mathcal{S}^{3}$ tokenizer in CosyVoice (Du et al., 2024a), a range of supervised semantic tokenizers have been proposed. For example, CosyVoice3 (Du et al., 2025) employs multi-task learning to improve prosody modeling in their tokenizer, while the Baichuan Audio Tokenizer (Li et al., 2025) introduced an additional mel reconstruction objective to capture both semantic and acoustic details. However, due to the lack of necessary acoustic information, the application of semantic tokens in generation tasks still requires additional generative models and reference audio to generate acoustic features. To fully leverage the semantic consistency inherent in semantic tokens while addressing their lack of acoustic information during generation tasks, we propose a method that preserves their semantic integrity by introducing a separate acoustic token stream.
@@
Acoustic tokens are derived from neural audio codecs, which compress continuous audio signals into discrete tokens while reconstructing high-fidelity waveforms (Niu et al., 2024). Speech codecs typically adopt the VQ-GAN paradigm (Esser et al., 2021), where a VQ-VAE (Van Den Oord et al., 2017)-based generator performs the encode–quantize–decode process, and adversarial discriminators distinguish real from synthetic speech to improve perceptual quality. SoundStream (Zeghidour et al., 2021) introduced residual vector quantization (RVQ) to improve quantization efficiency, while Encodec (Défossez et al., 2022) incorporated LSTMs into the encoder–decoder structure to enhance compression. More recently, several single-codebook codecs have emerged, including Single-Codec (Li et al., 2024), WavTokenizer (Ji et al., 2024b), and BigCodec (Xin et al., 2024), which replace multi-layer quantizers with a single codebook. These designs achieve ultra-low-bitrate compression while simplifying and accelerating downstream modeling. However, the absence of explicit semantic consistency constraints limits their effectiveness in recognition-oriented tasks and hinders compatibility with text-based LMs (Wang et al., 2025a).
@@
To enhance semantic representation, recent speech codecs have introduced semantic features through distillation or injection. For instance, SpeechTokenizer (Zhang et al., 2023) leverages HuBERT (Hsu et al., 2021) to guide the first RVQ layer toward encoding semantic information, while Mimi (Défossez et al., 2024) distills WavLM (Chen et al., 2022a) features into a separate VQ module. X-Codec (Ye et al., 2025a, b) introduces semantic injection via an auxiliary semantic module, and XY-Tokenizer (Gong et al., 2025) further enhances textual alignment through an LLM-based ASR objective. However, these methods rely on fused semantic–acoustic tokens that must jointly support semantic prediction and spectrogram reconstruction, limiting both semantic fidelity and reconstruction quality. Although SemantiCodec (Liu et al., 2024), built on AudioMAE (Huang et al., 2022), seeks to disentangle semantic and acoustic streams, our analysis indicates that the separation remains incomplete. In contrast, SAC employs a dual-stream semantic–acoustic design, enabling both stronger semantic representation and improved reconstruction.
@@ 3.1 Model Architecture
SAC is built on the VQ-GAN framework (Esser et al., 2021), which follows a VQ-VAE architecture (Gârbacea et al., 2019) to reconstruct raw speech. As shown in Fig. 2, SAC comprises a dual-stream encoder–quantizer and a unified codec decoder, complemented by auxiliary modules for semantic and speaker feature supervision. The details of each component are described in the following subsections.
@@
To ensure that the semantic stream of SAC maintains strong semantic consistency, we employ a pre-trained semantic tokenizer, keeping its parameters frozen throughout SAC training. Specifically, we use the speech tokenizer proposed in Zeng et al. (2024b), which tokenizes input speech into discrete tokens at a frame rate of 12.5 Hz. Formally, given an input waveform $x$, the semantic tokenizer first extracts fine-grained continuous representations $\mathbf{S}_{c}$ at 50 Hz, which also serve as the target for auxiliary semantic supervision. A temporal pooling layer then downsamples $\mathbf{S}_{c}$ to 12.5 Hz, producing $\mathbf{S}$. These features are quantized via a vector quantization layer to obtain discrete semantic tokens and their corresponding quantized embeddings $\mathbf{S}_{q}$. To achieve temporal alignment with the acoustic embeddings, $\mathbf{S}_{q}$ is further upsampled by a ConvNeXt-based adapter (Liu et al., 2022), resulting in the semantic features $\mathbf{S}_{q}^{\prime}$.
@@
The acoustic token serves to complement the acoustic details missing from the semantic token. Specifically, we adopt the Encodec architecture (Défossez et al., 2022), which employs stacked convolutional and temporal downsampling layers with stride $\tau$ to extract frame-level acoustic representations $\mathbf{A}$. Following DAC (Kumar et al., 2023), we apply factorized code projections to map $\mathbf{A}$ into a lower-dimensional embedding space and perform single-codebook quantization based on $L_{2}$ distances. To mitigate codebook underutilization, entries that remain inactive over prolonged training intervals are reinitialized with randomly sampled embeddings drawn from the current batch (Dhariwal et al., 2020). Since acoustic information is inherently more fine-grained than semantic information, we adopt higher frame rates for its representation, specifically 25 Hz for low-bitrate settings and 50 Hz for high-bitrate settings. The corresponding strides $\tau$ are set to $(2,2,4,5,8)$ and $(2,4,5,8)$, respectively, yielding temporal reduction factors of 640 and 320 for input audio sampled at 16 kHz.
@@
The quantized acoustic embeddings $\mathbf{A}_{q}$ are concatenated with the semantic embeddings $\mathbf{S}_{q}^{\prime}$ along the feature dimension to form a unified representation $\mathbf{U}$. This joint representation is then processed by a ConvNeXt-based prenet, which upsamples it to 50 Hz, producing the fused feature sequence $\mathbf{F}$. The fused representation $\mathbf{F}$ integrates linguistic information from the semantic stream with timbre and acoustic detail from the acoustic stream. Subsequently, $\mathbf{F}$ is passed through a mirrored decoder composed of stacked convolutional and temporal upsampling layers to reconstruct the waveform $\tilde{x}$, with deconvolution strides set to $\tau=(8,5,4,2)$. Following the design of “X-shaped” codec models (Ye et al., 2025a), we introduce an auxiliary semantic reconstruction objective to ensure that key linguistic information is preserved during decoding. Specifically, $\mathbf{F}$ is fed into a CNN-based semantic decoder to predict the reconstructed semantic features $\tilde{\mathbf{S}}_{c}$. A mean squared error (MSE) loss,
@@ 3.2 Auxiliary Speaker Feature Supervision
While the acoustic stream effectively captures fine-grained spectral details, it may insufficiently model global timbre characteristics. To mitigate this limitation and improve timbre reconstruction, we introduce explicit speaker feature supervision. To be specific, we employ ERes2Net (Chen et al., 2023) to extract speaker embeddings $\mathbf{S}_{p}$ as the supervision target. For prediction, we compute the temporal mean and variance of the fused representation $\mathbf{F}$ and concatenate them into a global feature $\mathbf{f}$. This vector is passed through a lightweight two-layer MLP projector to generate the predicted speaker embedding $\tilde{\mathbf{S}}_{p}$. An MSE loss is then applied between $\tilde{\mathbf{S}}_{p}$ and $\mathbf{S}_{p}$ to encourage accurate modeling of timbre information:
@@ 3.3 Training Objectives
For the acoustic stream, the codebook is optimized by minimizing the $L_{2}$ distance between the encoder outputs and their quantized embeddings, with gradients propagated using the straight-through estimator (STE) (Bengio et al., 2013). The VQ loss $\mathcal{L}_{\text{vq}}$ also incorporates a commitment term that constrains encoder outputs to remain close to their assigned codebook entries.
@@
We employ a multi-period discriminator (MPD) (Kong et al., 2020) and a multi-scale STFT-based discriminator (MS-STFT) (Défossez et al., 2022), following Xin et al. (2024). The discriminators are optimized using the least-squares GAN objective (Mao et al., 2017). For the generator, we apply both an adversarial loss $\mathcal{L}_{\text{adv}}$ and a feature matching loss $\mathcal{L}_{\text{feat}}$, the latter computed as the $L_{1}$ distance between intermediate feature maps of real and generated audio.

### v2

@@ Abstract
Speech codecs that convert continuous speech signals into discrete tokens have become essential for speech language models. However, existing codecs struggle to balance high-quality reconstruction with semantically rich representations, limiting their effectiveness in both generative and understanding tasks. In this work, we propose SAC, a neural speech codec with semantic-acoustic dual-stream quantization. By disentangling semantic and acoustic modeling into two dedicated streams, SAC enables each to be optimized for its respective role. Comprehensive evaluations show that SAC achieves strong reconstruction performance across diverse bitrates under both clean and noisy conditions, with particularly high scores on UTMOS and WER, indicating superior naturalness and intelligibility. Moreover, SAC substantially surpasses prior codecs in semantic representation, approaching the level of continuous self-supervised embeddings. When used as a tokenizer for LLM-based text-to-speech, SAC enables a single-stage autoregressive (AR) TTS model that clearly outperforms state-of-the-art AR systems. Our disentanglement analysis further validates the effectiveness of the dual-stream design, offering new potential for controllable speech generation. The code and pre-trained models are available at https://github.com/Soul-AILab/SAC.  [^1] Demo at https://sac-codec.github.io.
@@ 1 Introduction
With the rapid advancement of large language models (LLMs), speech language models (SLMs) have emerged by extending text-based LLMs with speech modalities (survey-pengjing; wavchat). Central to these models is the speech tokenizer, which discretizes continuous speech waveforms into token sequences, thereby enabling seamless integration with token-based language models (review-yiwei). Leveraging such speech tokens, SLMs have driven progress in a wide range of downstream applications, including text-to-speech (TTS) (fireredtts2; spark-tts), speech understanding (token-understanding), and spoken dialogue systems  (Qwen3-Omni; glm-4-voice; slam-omni; moshi).
Figure 1: Comparison of codecs on speech reconstruction. The x-axis shows WER, reflecting speech intelligibility, while the y-axis presents UTMOS, reflecting objective naturalness. Circle size indicates bitrate.
Semantic tokens are among the most widely used tokens in speech processing. These tokens are typically derived from either self-supervised models (case-study-yifan) or supervised models (Cosyvoice; Cosyvoice2), making them effective in capturing semantic meaning. However, the absence of essential acoustic information significantly limits the applicability of semantic tokens. In contrast, acoustic tokens are usually generated by neural audio codecs trained with reconstruction objectives (encodec; Soundstream). While this approach preserves fine-grained acoustic details, the lack of semantic supervision results in weaker alignment with semantic content and reduced compatibility with text-based LMs (AudioCodecBench; CodecBench).
To enhance semantic representation, recent advancements in speech codecs have explored incorporating semantic supervision during training. For example, SpeechTokenizer (speechtokenizer) employs semantic distillation, where representations from an SSL model guide the output of the first residual vector quantization (RVQ) layer. The X-Codec series (xcodec; llasa-xcodec2) and XY-Tokenizer (XY-Tokenizer) adopt an “X-shaped” paradigm, explicitly injecting pre-trained semantic features and fusing them with acoustic embeddings before quantization. However, while these approaches improve semantic alignment compared to methods without semantic constraints, they still fall short of pure semantic tokens in terms of semantic relevance. This raises a central question: Can semantic and acoustic tokens be disentangled at the token level, allowing each to specialize in its respective role? In this work, we explore such a dual-stream design that decouples semantic and acoustic modeling into independent streams, and show that it leads to improved semantic representation and reconstruction performance.
@@
Acoustic Dual-Stream Neural Speech Codec. Unlike prior approaches that inject semantic supervision into codecs, SAC complements the semantic tokens by introducing a separate acoustic token stream, which provides the essential acoustic information missing from the semantic tokens, all while ensuring the integrity of the semantic representations. Specifically, in the semantic stream, we adopt a pre-trained speech tokenizer (glm4voice-tokenizer) to extract semantic tokens aligned with linguistic content, keeping it frozen during training to ensure faithful retention of semantic information. In the acoustic stream, we follow the design of neural audio codecs (dac), employing temporally distributed acoustic tokens to capture the essential acoustic information, e.g., timbre and emotional attributes, that is missing from the semantic tokens. This dual-stream design unifies the complementary strengths of both within a single framework: semantic tokenizers excel in speech understanding and dialogue tasks (glm-4-voice; kimi-audio), while acoustic tokenizers are particularly effective in generative modeling (valle; valle2). At the decoding stage, a ConvNeXt-based (convnext) prenet is employed to fuse the two streams of embeddings, followed by a codec decoder that reconstructs the waveform. Experiments show that SAC delivers strong reconstruction and semantic relevance, while also supporting competitive performance in downstream LLM-based TTS tasks. Our main contributions can be summarized as follows:
  • We propose SAC, a semantic–acoustic dual-stream neural speech codec. By explicitly disentangling speech encoding into parallel semantic and acoustic streams, SAC enables each pathway to specialize in modeling linguistic content and acoustic detail, respectively.
  • SAC outperforms existing codecs with strong speech reconstruction quality and semantic representation across different bitrates.
  • We validate SAC in TTS downstream tasks, where SAC-based single-stage AR models significantly surpass SOTA pure AR systems in both intelligibility and objective naturalness.
@@ 2 Related Work
Semantic tokens are derived from quantized representations of self-supervised learning (SSL) models or supervised models. They are strongly correlated with textual semantics but are not primarily designed for waveform reconstruction. Following the introduction of the $\mathcal{S}^{3}$ tokenizer in CosyVoice (Cosyvoice), a range of supervised semantic tokenizers have been proposed. For example, CosyVoice3 (Cosyvoice3) employs multi-task learning to improve prosody modeling in their tokenizer, while the Baichuan Audio Tokenizer (Baichuan-audio) introduced an additional mel reconstruction objective to capture both semantic and acoustic details. However, due to the lack of necessary acoustic information, the application of semantic tokens in generation tasks still requires additional generative models and reference audio to generate acoustic features. To fully leverage the semantic consistency inherent in semantic tokens while addressing their lack of acoustic information during generation tasks, we propose a method that preserves their semantic integrity by introducing a separate acoustic token stream.
@@
Acoustic tokens are derived from neural audio codecs, which compress continuous audio signals into discrete tokens while reconstructing high-fidelity waveforms (NDVQ). Speech codecs typically adopt the VQ-GAN paradigm (vqgan), where a VQ-VAE (vqvae)-based generator performs the encode–quantize–decode process, and adversarial discriminators distinguish real from synthetic speech to improve perceptual quality. SoundStream (Soundstream) introduced residual vector quantization (RVQ) to improve quantization efficiency, while Encodec (encodec) incorporated LSTMs into the encoder–decoder structure to enhance compression. More recently, several single-codebook codecs have emerged, including Single-Codec (Single-codec), WavTokenizer (wavtokenizer), and BigCodec (bigcodec), which replace multi-layer quantizers with a single codebook. These designs achieve ultra-low-bitrate compression while simplifying and accelerating downstream modeling. However, the absence of explicit semantic consistency constraints limits their effectiveness in recognition-oriented tasks and hinders compatibility with text-based LMs (AudioCodecBench).
@@
To enhance semantic representation, recent speech codecs have introduced semantic features through distillation or injection. For instance, SpeechTokenizer (speechtokenizer) leverages HuBERT (hubert) to guide the first RVQ layer toward encoding semantic information, while Mimi (moshi) distills WavLM (wavlm) features into a separate VQ module. X-Codec (xcodec; llasa-xcodec2) introduces semantic injection via an auxiliary semantic module, and XY-Tokenizer (XY-Tokenizer) further enhances textual alignment through an LLM-based ASR objective. However, these methods rely on fused semantic–acoustic tokens that must jointly support semantic prediction and spectrogram reconstruction, limiting both semantic fidelity and reconstruction quality. Although SemantiCodec (semanticodec), built on AudioMAE (audiomae), seeks to disentangle semantic and acoustic streams, our analysis indicates that the separation remains incomplete. In contrast, SAC employs a dual-stream semantic–acoustic design, enabling both stronger semantic representation and improved reconstruction.
@@ 3.1 Model Architecture
SAC is built on the VQ-GAN framework (vqgan), which follows a VQ-VAE architecture (vqvae-speech) to reconstruct raw speech. As shown in Fig. 2, SAC comprises a dual-stream encoder–quantizer and a unified codec decoder, complemented by auxiliary modules for semantic and speaker feature supervision. The details of each component are described in the following subsections.
@@
To ensure that the semantic stream of SAC maintains strong semantic consistency, we employ a pre-trained semantic tokenizer, keeping its parameters frozen throughout SAC training. Specifically, we use the speech tokenizer proposed in glm4voice-tokenizer, which tokenizes input speech into discrete tokens at a frame rate of 12.5 Hz. Formally, given an input waveform $x$, the semantic tokenizer first extracts fine-grained continuous representations $\mathbf{S}_{c}$ at 50 Hz, which also serve as the target for auxiliary semantic supervision. A temporal pooling layer then downsamples $\mathbf{S}_{c}$ to 12.5 Hz, producing $\mathbf{S}$. These features are quantized via a vector quantization layer to obtain discrete semantic tokens and their corresponding quantized embeddings $\mathbf{S}_{q}$. During training, the semantic tokenizer is kept frozen to ensure that the semantic stream focuses exclusively on linguistic content without being biased toward acoustic details. To achieve temporal alignment with the acoustic embeddings, $\mathbf{S}_{q}$ is further upsampled by a ConvNeXt-based adapter (convnext), resulting in the semantic features $\mathbf{S}_{q}^{\prime}$.
@@
The acoustic token serves to complement the acoustic details missing from the semantic token. Specifically, we adopt the Encodec architecture (encodec), which employs stacked convolutional and temporal downsampling layers with stride $\tau$ to extract frame-level acoustic representations $\mathbf{A}$. Following DAC (dac), we apply factorized code projections to map $\mathbf{A}$ into a lower-dimensional embedding space and perform single-codebook quantization based on $L_{2}$ distances. To mitigate codebook underutilization, entries that remain inactive over prolonged training intervals are reinitialized with randomly sampled embeddings drawn from the current batch (Jukebox). Since acoustic information is inherently more fine-grained than semantic information, we adopt higher frame rates for its representation, specifically 25 Hz for low-bitrate settings and 50 Hz for high-bitrate settings. The corresponding strides $\tau$ are set to $(2,2,4,5,8)$ and $(2,4,5,8)$, respectively, yielding temporal reduction factors of 640 and 320 for input audio sampled at 16 kHz.
@@
The quantized acoustic embeddings $\mathbf{A}_{q}$ are concatenated with the semantic embeddings $\mathbf{S}_{q}^{\prime}$ along the feature dimension to form a unified representation $\mathbf{U}$. This joint representation is then processed by a ConvNeXt-based prenet, which upsamples it to 50 Hz, producing the fused feature sequence $\mathbf{F}$. The fused representation $\mathbf{F}$ integrates linguistic information from the semantic stream with timbre and acoustic detail from the acoustic stream. Subsequently, $\mathbf{F}$ is passed through a mirrored decoder composed of stacked convolutional and temporal upsampling layers to reconstruct the waveform $\tilde{x}$, with deconvolution strides set to $\tau=(8,5,4,2)$. Following the design of “X-shaped” codec models (xcodec), we introduce an auxiliary semantic reconstruction objective to ensure that key linguistic information is preserved during decoding. Specifically, $\mathbf{F}$ is fed into a CNN-based semantic decoder to predict the reconstructed semantic features $\tilde{\mathbf{S}}_{c}$. A mean squared error (MSE) loss,
@@ 3.2 Auxiliary Speaker Feature Supervision
While the acoustic stream effectively captures fine-grained spectral details, it may insufficiently model global timbre characteristics. To mitigate this limitation and improve timbre reconstruction, we introduce explicit speaker feature supervision. To be specific, we employ ERes2Net (eres2net-sv) to extract speaker embeddings $\mathbf{S}_{p}$ as the supervision target. For prediction, we compute the temporal mean and variance of the fused representation $\mathbf{F}$ and concatenate them into a global feature $\mathbf{f}$. This vector is passed through a lightweight two-layer MLP projector to generate the predicted speaker embedding $\tilde{\mathbf{S}}_{p}$. An MSE loss is then applied between $\tilde{\mathbf{S}}_{p}$ and $\mathbf{S}_{p}$ to encourage accurate modeling of timbre information:
@@ 3.3 Training Objectives
For the acoustic stream, the codebook is optimized by minimizing the $L_{2}$ distance between the encoder outputs and their quantized embeddings, with gradients propagated using the straight-through estimator (STE) (ste). The VQ loss $\mathcal{L}_{\text{vq}}$ also incorporates a commitment term that constrains encoder outputs to remain close to their assigned codebook entries.
@@
We employ a multi-period discriminator (MPD) (hifi-gan) and a multi-scale STFT-based discriminator (MS-STFT) (encodec), following bigcodec. The discriminators are optimized using the least-squares GAN objective (lsgan). For the generator, we apply both an adversarial loss $\mathcal{L}_{\text{adv}}$ and a feature matching loss $\mathcal{L}_{\text{feat}}$, the latter computed as the $L_{1}$ distance between intermediate feature maps of real and generated audio.